CoolRekterizer:

What is the equation, in point-slope form, for a line that goes through ​ (2, −6)(2, −6) ​ and has a slope of ​ −34−34 ​ ? y−6=34(x−2)y−6=34(x−2) y−6=−34(x+2)y−6=−34(x+2) y+6=−34(x−2)y+6=−34(x−2) y+6=−34(x+2)

CoolRekterizer:

I suck at slopes but my decision is D, right?

Shadow:

\[y - y_{1} = m(x - x_{1})\] \[y + 6 = -34(x - 2)\]

Shadow:

I am assuming some of these answers (text) have been oddly resposted, like in the question.

CoolRekterizer:

FLVS

CoolRekterizer:

thanks.